OpenStudy (anonymous):

Determine the equation of th tangent line to the function f(x)=x/e^2x at the point (1,1/e^2)

OpenStudy (anonymous):

take the derivative of f(x) and plug 1 into your derivative to find the slope of your tangent line at that point

OpenStudy (anonymous):

but how do I derive (x/e^2x)?? thats where I am stuck

OpenStudy (anonymous):

wait isnt it using the quotient rule so the denominator should be e^2x squared right??

OpenStudy (experimentx):

do you know how to differentiate \[ f(x)=\frac{x}{e^{2x}} \]

OpenStudy (anonymous):

quotient rule right ??

OpenStudy (anonymous):

1* e^2x+x*2e^2x/e^2x squared??

OpenStudy (experimentx):

yep ... now put, x=1, what do you get?

OpenStudy (experimentx):

i mean put x=1 on your derivative.

OpenStudy (anonymous):

okay i need help beacuase i would be guessing \

OpenStudy (experimentx):

Dang!! just replace x=1 on 1* e^2x+x*2e^2x/e^2x squared??

OpenStudy (anonymous):

i get e^2 =2e^2/(e^2)^2

OpenStudy (anonymous):

can you help me experiment,please?

OpenStudy (experimentx):

?? no ...

OpenStudy (experimentx):

you made some mistake somewhere wolf says this is zero. http://www.wolframalpha.com/input/?i=D%5Bx%2Fe%5Ex%2C+x%5D%2F.x-%3E1

OpenStudy (experimentx):

also you messed up your derivative somewhere http://www.wolframalpha.com/input/?i=D%5Bx%2Fe%5Ex%2C+x%5D

OpenStudy (anonymous):

i give up

OpenStudy (anonymous):

i got -1/e^2 after plugging 1 into the derivative

OpenStudy (anonymous):

travis can you do it step by step if its not too much.

OpenStudy (experimentx):

do your derivative again ... wolf says the derivative is \[ \frac{1}{e^x} - \frac{x}{e^{x}}\] you probably messed up somewhere.

OpenStudy (experimentx):

Woops!! my fault ... i missed e^2x http://www.wolframalpha.com/input/?i=D%5Bx%2Fe%5E%282x%29%2C+x%5D%2F.x-%3E1

OpenStudy (experimentx):

now use this formula \[ y - y' = f'(x') (x-x')\] replace \((x', y') \to (1, 1/e) \) and replace your \( f'(1) \) by above.

OpenStudy (anonymous):

i used the quotient rule f'(x)=(e^2x-x2e^2x)/(e^2x)^2

OpenStudy (anonymous):

plug in 1 and you get (e^2-2e^2)/e^4

OpenStudy (anonymous):

this simplifies to (-e^2)/(e^4) which simplifies to -1/e^2